How many mL of a 0.184 M NaNO_3 solution contains How many mL of M ?
Answer #1

M = n /V

M = molarity = 0.184

n = number of moles = 0.113

V = volume in L

V = n /M

V = 0.113 / 0.184

V = 0.614 L = 614 mL

answer = option C = 614 mL